OpenStudy (anonymous):

what are the different ways to rewrite 2cosx

OpenStudy (anonymous):

Do you mean cos(2x) ? There are 3 different ways to write that.

OpenStudy (anonymous):

No, 2cosx

OpenStudy (anonymous):

I am trying to solve the equation\[tanx +secx =2cosx\]

OpenStudy (anonymous):

you see how you can get the left side to be (sin x + 1) / cos x ?

OpenStudy (anonymous):

yeah I did that but did not know where to go from there

OpenStudy (anonymous):

Now multiply both sides by cos x and you get sin x + 1 = (cos x)^2

OpenStudy (anonymous):

now rewrite (cosx)^2 as 1 - (sinx)^2

OpenStudy (anonymous):

oops i messed up ... the right side will be 2(cos x)^2

OpenStudy (anonymous):

yeah I was trying to type that

OpenStudy (anonymous):

you are able to get those equations written a lot faster than I can

OpenStudy (anonymous):

that's why I don't know what to do with the 2 on the outside

OpenStudy (anonymous):

You'll get a quadratic equation with only sin's. Get everything on one side and 0 on the other and solve

OpenStudy (anonymous):

how on earth do you use the quadratic formula with trig functions?

OpenStudy (anonymous):

I got it to \[\sin ^{2}x+sinx-1=0\] but then what?

OpenStudy (anonymous):

I get\[2\sin ^{2}x+sinx-1=0\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

oh you are right, I didn't distribute correctly

OpenStudy (anonymous):

\[sinx+1=2\cos^{2}x\]\[sinx+1=2(1-\sin ^{2}x)\] Distribute the 2 over the 1 and - sin^2x. Need i go further

OpenStudy (anonymous):

ohh. ok cool

OpenStudy (anonymous):

it'll factor into(2sinx-1)(sinx +1)=0

OpenStudy (anonymous):

oh perfect. Thanks! I am new to this site and would love to give you a medal, but I don't know how to do that either

OpenStudy (anonymous):

me too. i'm just really bored. I need to figure out a way to get paid to do this LOL!

OpenStudy (anonymous):

No kidding! That would be awesome. you can make some good money tutoring, but this would be way easier